# Recent Advances in Discrete Speech Tokens: A Review

### v1

@@ III-D1 Motivation
Another line of mixed-objective acoustic tokens is disentanglement. A prominent research direction is the disentanglement of speaker timbre information, as this is a global trait among all the speech information aspects. Encoding speaker information into every token timestep is redundant; thus, removing the global speaker timbre can make the information in acoustic tokens more compact and reduce the necessary bitrate. Speaker-decoupled speech tokens can alleviate the modeling burden for downstream tasks. For example, a TTS model using these tokens can achieve independent control over prosody and speaker identity.
The disentanglement of speaker timbre also enables an acoustic token to perform voice conversion (VC), as timbre from the target speaker can be easily combined with the speaker-agnostic content tokens from the source speech. Note that in Section III, it is mentioned that some codecs introduce a global encoder or low-rate prosody encoder to reduce the necessary bitrate of time-variant tokens [20, 21, 71, 81]. They have already demonstrated some ability to decouple global speaker timbre and local contents, albeit in an implicit manner through the natural information bottleneck in the VQ layers. In this section, we elaborate on explicit methods, which typically involve specialized training techniques and criteria to achieve disentanglement.
@@ III-D2 Approaches
SSVC [26] is one of the pioneering efforts in this direction, which is the basis for BASE-TTS [112]. SSVC designs two coupled regressors from WavLM: a speaker regressor and a content regressor. These regressors are essentially attention modules on every WavLM layer. The speaker regressor is used to train a speaker extractor by contrastive loss to produce discriminative speaker embeddings, similar to the GE2E loss [113] in speaker verification. The output from the content regressor is quantized by RVQ before being concatenated with speaker embeddings and reconstructed into waveforms by a BigVGAN [114] vocoder. Disentanglement is mostly ensured by a GRL on a cosine distance loss between the speaker extractor outputs from the speaker regressor and content regressor. Similarly, PromptCodec minimizes an SSIM loss [115] between content and timbre representations, with the help of a pretrained speaker verification model.
Such GRL technique is not limited to disentangling speaker timbre alone. FACodec [27] employs supervised decoupling to factorize speech into speaker timbre, content, prosody, and acoustic detail information. The timbre extractor in FACodec is optimized via a speaker classification loss. For the other components – prosody, content, and acoustic detail – separate RVQ modules are applied prior to the supervised decoupling process. In the content branch, frame-aligned phonemes are predicted, and GRL is applied to the normalized F0. In the prosody branch, normalized F0 is predicted, and GRL is applied to the frame-aligned phonemes. Meanwhile, in the acoustic detail branch, GRL is performed using both phonemes and F0. These three quantized features are then combined before applying GRL with the speaker classification information. Finally, the decoder of FACodec integrates all four information branches to reconstruct the speech signal.
@@
For speaker disentanglement, a more straightforward approach is to apply speaker timbre perturbations to speech signals and leverage the strong information bottleneck created by the discrete VQ module. When the encoder is unable to learn sufficient timbre information, and the decoder is provided with prominent timbre, the bottleneck in the middle will naturally prevent timbre from being encoded [116]. Such an idea is adopted in LSCodec [28] to achieve speaker decoupling and ultra-low-bitrate coding. LSCodec leverages continuous WavLM features to represent speaker timbre for its remarkable speaker verification ability [117, 118]. These features are fed to a Conformer-based decoder by position-agnostic cross attention [119, 120]. A stretch-based speaker perturbation algorithm is applied to the input waveform to facilitate speaker disentanglement. The training process of LSCodec involves two stages: first, a speech VAE is trained obtain a preliminary speaker-decoupled continuous space. Subsequently, this continuous space is discretized into a VQ-VAE. Through this speaker disentanglement approach, LSCodec achieves high-quality speech reconstruction and voice conversion (VC) using only a single codebook, with an extremely low bitrate of 0.25 kbps.
@@
Apart from the disentanglement of speaker timbre, source separation has also been explored in the context of acoustic tokens. SD-Codec [29] proposes to decouple different audio sources in the neural codec, like speech, music, and sound effects. By employing a shared encoder and decoder architecture with multiple parallel RVQ modules, SD-Codec aims to separate different audio sources into distinct RVQ streams. This approach allows for more efficient and targeted processing of each audio component.
@@ IV-A2 Approaches
This type of speech SSL models aims to learn representations by distinguishing a target sample (positives) from distractors (negatives) given an anchor [8]. They minimize the latent space similarity of negative sample pairs and maximize that of the positive pairs. For semantic tokens, vq-wav2vec [30] and wav2vec 2.0 [31] are two representative contrastive SSL models. They both involve a quantizer to produce localized features that is contrastively compared to contextualized continuous features. Vq-wav2vec [30] uses pure CNN blocks while wav2vec 2.0 [31] adopts a Transformer for stronger contextual features. Both use GVQ quantizers with two groups to expand the VQ space. Wav2vec 2.0 has also been extended to massively multilingual versions [146, 134, 147].
@@
This type of speech SSL models incorporates an external target for prediction, either from signal processing features or another teacher network. A popular line of work is hidden unit BERT (HuBERT) [32]. It takes raw waveforms as inputs, applies random masks on the hidden representations before Transformer contextual blocks, and then predicts quantized MFCC targets. It can take more self-iterations by using a trained HuBERT teacher model and applying k-means clustering as targets. WavLM [33] augments HuBERT by additional speaker and noise perturbations to achieve superior performance in more paralinguistic-related tasks. There are no inner quantizers in both models, so external quantization like k-means clustering is necessary to obtain semantic tokens. BEST-RQ [144] changes the prediction target to the output of a random projection quantizer. Data2vec [148, 149] proposes a general Transformer-based teacher-student masked prediction framework the masked and original view of data are fed to the student and teacher respectively, and the student network predicts the teacher outputs. The next-token prediction criterion from language models (LMs) have also been adopted into speech SSL [150, 151], either with or without a pretrained text LM. This method emphasizes the autoregressive prediction property of learned tokens that may be better suited for the LM use case.
@@ IV-B1 Motivation
As SSL tokens feature semantic or phonetic information, a major concern is to improve the resistance against perturbations in the input signal. This kind of invariance includes noise and speaker aspects that don’t affect the contents of speech. Noise invariance refers to the property that semantic tokens should be invariant against signal augmentations such as additive noise, reverberations, etc. Speaker invariance is similar to speaker-disentangled acoustic tokens, where speaker information needs to be removed in the semantic tokens. In the training process, perturbations are often explicitly introduced in these perturbation-invariant SSL models. The original and perturbed view of an utterance are both fed to the same network (or teacher and student networks), and an external loss to reduce the impact of perturbation is applied. The middle part of Fig.7 depicts a typical perturbation-invariant SSL model.
@@ IV-B2 Approaches
The perturbations can either be designed to augment the acoustics or alter the speaker timbre, depending on the objective of invariance. These perturbations usually preserve temporal alignments, meaning that the perturbed utterance and the original one are strictly synchronized. For noise-invariant SSL tokens, basic signal variations like time stretching, pitch shifting, additive noise, random replacing and reverberation are commonly applied [38, 154, 37]. SpecAugment [155] is also used in [35]. Typical speaker timbre perturbations include formant and pitch scaling as well as random equalization [34, 36, 156]. In contrast, random time stretching is applied as speaker perturbation in [157], which alters the tempo in each random segment.
@@
Another method to achieve invariance is to minimize some distance metrics between the representations extracted from the original and perturbed utterances. In existing perturbation-invariant SSL models, this is typically accomplished using a cross-entropy loss between the underlying distributions in the VQ module of the SSL model. NAST [37] trains a Gumbel-based VQ-VAE on HuBERT features and enforces similarity between the Gumbel distributions Eq.(3) derived from the original and augmented utterances. Spin [36] and DC-Spin [156] explore a speaker-invariant clustering algorithm for HuBERT features. Similar to NAST [37], Spin employs a cross-entropy loss to ensure that the distributions over codebook entries are similar between the original and perturbed utterances. Spin uses a distribution smoothing technique before pushing the distributions to be similar, thereby preventing collapse into a trivial solution [36]. This distribution-based approach forces the same content to be quantized to the same index regardless of its acoustic conditions. DC-Spin [156] uses Spin units to train a HuBERT model and extends the Spin algorithm to incorporate two VQ codebooks, both optimized with the same objective The auxiliary codebook is designed to be larger than the primary one, allowing for more fine-grained acoustic details Additionally, DC-SPIN explores fine-tuning with mel reconstruction and supervised ASR, which are anticipated to further enhance speaker invariance.
@@ IV-C Semantic Tokens from Supervised Models
As representing semantic or phonetic information is the major purpose of semantic tokens, a more direct way to achieve this is through supervised learning. A famous example shown at the bottom of Fig.7 is the $\mathcal{S}^{3}$ Tokenizer from CosyVoice [40]. It places a single-codebook VQ layer between two Transformer encoder modules and optimizes the network through an ASR next token prediction loss, similar to Whisper [39]. The same method is adopted in [161, 163] where the frame rate is further reduced to 12.5Hz. CosyVoice 2 [162] improves $\mathcal{S}^{3}$ Tokenizer by replacing plain VQ with FSQ for better codebook utilization. Note that in this kind of supervised semantic tokens, it is the output of the VQ layer that serves as tokens. This allows for more preservation of paralinguistic information than directly transcribing speech into text. These supervised tokenizers are trained on massive paired speech-text data, and have demonstrated rich speech content understanding capabilities [40, 164].
@@ IV-D Length Reduction by Deduplication and Acoustic BPE
In most cases, discrete speech tokens have a frame rate ranging from 25 to 100Hz, as can be found in Table I and Table II. This leads to a large discrepancy in lengths between speech representations and the underlying text modality. Most token sequences are tens of times longer than their corresponding phoneme sequences, not to mention the inner semantics. This discrepancy has been a critical issue in building decoder-only TTS and other LM-based speech generation tasks, since longer sequences result in harder training and more unstable inference. Therefore, length reduction techniques have been proposed to address this issue. These methods are inspired by language processing techniques and are thus more closely related to semantic tokens. It is important to note that, although these length reduction methods are universal across token types, they are less frequently applied in the context of acoustic tokens. This is because acoustic tokens usually involve multiple VQ streams that complicate the token-level operations.
@@
A common approach to reduce token sequence lengths is deduplication [135, 165], i.e. removing the repeated consecutive tokens in a sequence. Since the encoded continuous features are often close in consecutive frames where the speech dynamics do not change rapidly , they are likely to be quantized to the same unit. Therefore, removing these redundant tokens can yield a more phonetic representation. Consider an original token stream $[a,a,a,b,b,a,c]$. After deduplication, the sequence becomes $[a,b,a,c]$, with corresponding durations $[3,2,1,1]$. In downstream generation tasks, the deduplicated tokens are used for generative modeling, and a unit-to-speech model (similar to TTS) is employed to convert the tokens back to acoustic signals [4].
@@
In addition to deduplication, another popular approach to reducing the length of speech token sequences is acoustic byte-pair encoding (BPE) [^6] The term “acoustic” here is used to distinguish it from traditional BPE applied to text tokens, rather than referring to “acoustic tokens”. or so-called subword modeling [166, 167, 135, 168, 169]. Similar to text BPE [170], acoustic BPE iteratively merges the two most frequent consecutive tokens and adds the merged token to the vocabulary. After training on a corpus, a deterministic BPE mapping is established between original token combinations and the new vocabulary. This mapping enables a lossless compression algorithm, allowing tokens to be perfectly reconstructed after BPE encoding and decoding. This operation identifies certain morphological patterns in token sequences and offers a powerful way to remove redundant tokens. The encoded BPE tokens are used for downstream generation tasks, and original tokens are recovered by the deterministic BPE mapping before converting to signals. The application of acoustic BPEs on HuBERT semantic tokens has been reported to yield a significant speed and performance gain in ASR [135, 165], spoken language modeling [168, 169] and TTS [171, 137].
@@
Although deduplication is a simple and training-free method, acoustic BPE offers unique advantages. Acoustic BPE can identify redundant patterns that are not simply repetitions. Also, deduplication discards the duration information of every token in the resulting sequence. This might be a concern for downstream tasks, as certain rhythmic information could reside in such token repetitions. In contrast, acoustic BPE preserves duration by encoding repetition in different lengths into distinct new tokens. Acoustic BPE is also flexible in terms of target vocabulary size, which can be adjusted based on the desired length reduction ratio and downstream performance.
@@
We visualize the length reduction effect of BPE on different speech tokens in Fig.8. In addition to semantic tokens from various models and different k-means clusters in HuBERT, we also experiment with acoustic tokens. For acoustic tokens with multiple codebooks, we apply BPE only to the first quantizer. This approach has the greatest potential in the current speech generation paradigm [6]. From Fig.8, it is evident that different types of tokens exhibit very distinct patterns. Semantic tokens generally show significant length reduction when applying BPE, especially for HuBERT models with fewer k-means clusters. For single-codebook acoustic tokens, speaker-decoupled LSCodec demonstrates more reduction, whereas WavTokenizer and BigCodec are almost unaffected by BPE. For a single RVQ layer among multiple-codebook acoustic tokens, the reduction effect is also significant. These findings suggest that the effect of BPE is negatively correlated with the information density in the speech tokens: the less information, the more length reduction achieved by BPE.
@@ IV-E Variable Frame Rate Tokens and Unit Discovery
VFR tokens are closely related to acoustic unit discovery. As speech lacks a natural boundary of phonetic units [8], there are much research efforts to find and locate the underlying acoustic units behind speech utterances in an unsupervised manner [173, 174, 175, 176]. This is particularly of interest for low-resource languages. These discovered units can guide the boundary segmentation of VFR tokens. To this end, VFR tokens can be interesting not only because it might reduce the necessary bitrate, but also it can introduce a strong inductive bias that linguistic knowledge is encoded [172].
A recent direction of VFR tokens is to discover acoustic units from an SSL model. Deduplicated tokens and acoustic BPE themselves can be regarded as VFR tokens. Sylber [177] and SyllableLM [44] take similar approaches that first locate acoustic boundaries from existing HuBERT models, and then train another HuBERT student with segment-level pooled targets between boundaries. The final HuBERT embeddings undergo the same segment-level pooling and kmeans clustering procedure to produce tokens at a very low frame rate ($\leq 5$Hz) that align well with syllables.
Boundary prediction can be involved to achieve frame rate variability in the training process, where a specific model predicts frame-level boundaries and is trained together with other network modules. Cuervo et al. [178] applies a hard boundary predictor into a contrastive speech SSL model for variable-rate downsampling and trains it using reinforcement learning. Hwang et al. [157] uses a soft predictor instead and performs downsampling through a soft pooling mechanism. Different from boundary prediction, a slowness constraint is imposed into a VQ-VAE in [172] that forces the latent features to vary slowly along time, after which run-length encoding (i.e. deduplication with duration saved) is applied on the quantized codes. However, these approaches are rarely adopted in the context of discrete speech tokens today, and there have barely been a VFR acoustic token till now.

### v2

@@ III-D1 Motivation
Another line of mixed-objective acoustic tokens is disentanglement. A prominent research direction is the disentanglement of speaker timbre information, as this is a global trait among all the speech information aspects. Encoding speaker information into every token timestep is redundant; thus, removing the global speaker timbre can make the information in acoustic tokens more compact and reduce the necessary bitrate. Speaker-decoupled speech tokens can alleviate the modeling burden for downstream tasks. For example, a TTS model using these tokens can achieve independent control over prosody and speaker identity. The disentanglement of speaker timbre also enables an acoustic token to perform voice conversion (VC), as timbre from the target speaker can be easily combined with the speaker-agnostic content tokens from the source speech.
Note that in Section III, it is mentioned that some codecs introduce a global encoder to reduce the necessary bitrate of time-variant tokens [21, 22, 72, 82]. They have already demonstrated some ability to decouple global speaker timbre and local contents, albeit in an implicit manner through the natural information bottleneck from VQ. In this section, we elaborate on explicit methods, which involve specialized training techniques and criteria to achieve disentanglement.
@@ III-D2 Approaches
SSVC [27] is one of the pioneering efforts in this direction. SSVC attempts to decouple content and speaker representations from WavLM features. The content branch is quantized via RVQ, and the speaker branch is trained using a contrastive loss to produce speaker embeddings. Disentanglement is enforced by a GRL between the speaker embeddings produced from the speaker branch and the content representations. Similarly, PromptCodec minimizes an SSIM loss [109] between content and speaker representations, with the help of a pretrained speaker verification model.
Such GRL technique is not limited to disentangling speaker timbre alone. FACodec [28] employs supervised decoupling to factorize speech into speaker timbre, content, prosody, and acoustic detail information. The timbre extractor in FACodec is optimized via a speaker classification loss. For the other components – prosody, content, and acoustic detail – separate RVQ modules are applied prior to the supervised decoupling process. For each component, some supervision signal with the desired information is applied, and GRL is employed to other non-related information components. These three quantized features are then combined before applying GRL with the speaker information. Finally, the decoder integrates all four information branches to reconstruct the speech signal.
@@
For speaker disentanglement, a more straightforward approach is to apply speaker timbre perturbations to speech signals and leverage the strong information bottleneck created by the discrete VQ module. When the encoder is unable to learn sufficient timbre information, and the decoder is provided with prominent timbre, the bottleneck in the middle will naturally prevent timbre from being encoded [110]. This idea is adopted in LSCodec [29] to achieve speaker decoupling and ultra-low-bitrate coding. LSCodec leverages continuous WavLM features to represent speaker timbre. These features are fed to a Conformer-based decoder by position-agnostic cross attention [111, 112]. A stretching-based speaker perturbation algorithm is applied to the input waveform to facilitate speaker disentanglement. The training process of LSCodec involves multiple stages where a VQ module is injected after constructing a speaker-decoupled continuous space. Through this approach, LSCodec achieves high-quality speech reconstruction and voice conversion using only a single codebook with very low bitrates.
@@
Apart from the disentanglement of speaker timbre, source separation has also been explored in the context of acoustic tokens. SD-Codec [30] proposes to decouple different audio sources in the neural codec, like speech, music, and sound effects, by employing multiple parallel RVQ modules. This approach allows for more efficient and targeted processing of each audio component.
@@ IV-A2 Approaches
This type of speech SSL models aims to learn representations by distinguishing a target sample (positives) from distractors (negatives) given an anchor [8]. They minimize the latent space similarity of negative pairs and maximize that of the positive pairs. For semantic tokens, vq-wav2vec [31] and wav2vec 2.0 [32] are two representative contrastive SSL models. They involve a quantizer to produce localized features that is contrastively compared to contextualized continuous features. Vq-wav2vec [31] uses pure CNN blocks while wav2vec 2.0 [32] adopts a Transformer for stronger capacity. Both use GVQ quantizers with two groups to expand the VQ space. Wav2vec 2.0 has also been extended to massively multilingual versions [139, 127, 140].
@@
This type of speech SSL models incorporates an external target for prediction, either from signal processing features or another teacher network. A popular line of work is HuBERT [33]. It takes raw waveforms as inputs, applies random masks on the hidden representations before Transformer contextual blocks, and then predicts k-means quantized targets from MFCC or another HuBERT teacher. WavLM [34] augments HuBERT by additional speaker and noise perturbations to achieve superior performance in more paralinguistic-related tasks. There are no inner quantizers in both models, so external quantization like k-means clustering is necessary to obtain semantic tokens. BEST-RQ [137] changes the prediction target to the output of a random projection quantizer. The next-token prediction criterion from language models (LMs) have also been adopted into speech SSL [141, 142], either with or without a pretrained text LM. This method emphasizes the autoregressive prediction property of learned tokens that may be better suited for the LM use case.
@@ IV-B1 Motivation
As SSL tokens feature semantic or phonetic information, a major concern is to improve the resistance against perturbations in the input signal. This kind of invariance includes noise and speaker aspects that don’t affect the contents of speech. Noise invariance refers to the invariance against signal augmentations such as additive noise, reverberations, etc. Speaker invariance aims to remove speaker information, similar to speaker-disentangled acoustic tokens. In the training process, perturbations are often explicitly introduced in these perturbation-invariant SSL models. The original and perturbed view of an utterance are both fed to the same network (or teacher and student networks), and an external loss to reduce the impact of perturbation is applied. The middle part of Fig.7 depicts a typical perturbation-invariant SSL model.
@@ IV-B2 Approaches
The perturbations can either be designed to augment the acoustics or alter the speaker timbre, depending on the objective of invariance. These perturbations usually preserve temporal alignments, meaning that the perturbed utterance and the original one are strictly synchronized. For noise-invariant SSL tokens, basic signal variations like time stretching, pitch shifting, additive noise, random replacing, reverberation, and SpecAugment [145] are commonly applied [39, 146, 38, 36]. Typical speaker timbre perturbations include formant and pitch scaling as well as random equalization [35, 37, 147]. In contrast, random time stretching is applied as speaker perturbation in [148], which alters the tempo in each random segment.
@@
Another method to achieve invariance is to minimize some distance metrics between the representations extracted from the original and perturbed utterances. In existing perturbation-invariant SSL models, this is typically accomplished using a cross-entropy loss between the underlying distributions in the VQ module of the SSL model. NAST [38] trains a Gumbel-based VQ-VAE on HuBERT features and enforces similarity between the Gumbel distributions Eq.(3) derived from the original and augmented utterances. Spin [37] and DC-Spin [147] explore a speaker-invariant clustering algorithm for HuBERT features. Similar to NAST [38], Spin employs a cross-entropy loss to ensure that the distributions over codebook entries are similar between the original and perturbed utterances. This distribution-based approach forces the same content to be quantized to the same index regardless of acoustic conditions.
@@ IV-C Semantic Tokens from Supervised Models
As representing semantic or phonetic information is the major purpose of semantic tokens, a more direct way to achieve this is through supervised learning. A famous example shown at the bottom of Fig.7 is the $\mathcal{S}^{3}$ Tokenizer from CosyVoice [41]. It places a single-codebook VQ layer between two Transformer encoder modules and optimizes the network through an ASR loss similar to Whisper [40]. The same method is adopted in [152, 154] where the frame rate is further reduced to 12.5Hz. CosyVoice 2 [153] improves $\mathcal{S}^{3}$ Tokenizer by replacing plain VQ with FSQ for better codebook utilization. Note that in this kind of supervised semantic tokens, it is the output of the VQ layer that serves as tokens. This allows for more preservation of paralinguistic information than directly transcribing speech into text. These supervised tokenizers are trained on massive paired speech-text data, and have demonstrated rich speech content understanding capabilities [41, 155].
@@ IV-D Length Reduction by Deduplication and Acoustic BPE
In most cases, the frame rate of discrete speech tokens ranges from 25 to 100Hz. This leads to a huge discrepancy in lengths between speech representations and the underlying text modality. This discrepancy has been a critical issue in building decoder-only TTS and other LM-based speech generation tasks, since longer sequences result in harder training and more unstable inference. Therefore, length reduction techniques have been proposed to address this issue. These methods are inspired by language processing techniques and are thus more closely related to semantic tokens. Note that although these length reduction methods are universal across token types, they are less frequently applied to acoustic tokens. This is because acoustic tokens usually involve multiple VQ streams that complicate token-level operations.
@@
A common approach to reduce token sequence lengths is deduplication [128, 156], i.e. removing the repeated consecutive tokens in a sequence. Since the encoded continuous features are often close in consecutive frames where the speech dynamics do not change rapidly, they are likely to be quantized to the same unit. Therefore, removing these redundant tokens can yield a more phonetic representation. When the deduplicated tokens are used for generative modeling, a unit-to-speech model (similar to TTS) should be employed to upsample the tokens and convert them back to acoustic signals [4].
@@
Another popular approach to reducing the length of speech token sequences is acoustic byte-pair encoding (BPE) [^4] The term “acoustic” here is used to distinguish it from traditional BPE applied to text tokens, rather than referring to “acoustic tokens”. or so-called subword modeling [157, 158, 128, 159, 160]. Similar to text BPE [161], acoustic BPE iteratively merges the two most frequent consecutive tokens and adds the merged token to the vocabulary. After training on a corpus, a deterministic BPE mapping is established between original token combinations and the new vocabulary. This mapping enables a lossless compression algorithm, allowing tokens to be perfectly reconstructed after BPE decoding. This operation can identify certain morphological patterns in token sequences, and offers a powerful way to remove redundant tokens. In practice, acoustic BPEs on HuBERT semantic tokens has demonstrated significant speed and performance gains in ASR [128, 156], spoken language modeling [159, 160] and TTS [162, 130].
@@
Although deduplication is a simple and training-free method, acoustic BPE offers several unique advantages over it. First, acoustic BPE can identify redundant patterns that are not simply repetitions, whereas deduplication only removes exact duplicates. Additionally, deduplication discards the duration information of every token in the resulting sequence. This could be problematic for downstream tasks, as important rhythmic information may reside in the repetitions of tokens. In contrast, acoustic BPE preserves duration information by encoding repetitions of varying lengths into distinct new tokens. Furthermore, acoustic BPE is flexible in terms of target vocabulary size, which can be adjusted based on the desired length reduction ratio and downstream performance.
@@
We visualize the length reduction effect of BPE on different speech tokens in Fig.8. In addition to semantic tokens from various models and different k-means clusters in HuBERT, we also experiment with acoustic tokens. For acoustic tokens with multiple codebooks, we apply BPE only to the first quantizer, in accordance with the current speech generation paradigm [6]. From Fig.8, it is evident that different types of tokens exhibit very distinct patterns. Semantic tokens generally show significant length reduction when applying BPE, especially for HuBERT models with fewer k-means clusters. For single-codebook acoustic tokens, speaker-decoupled LSCodec tokens shows more reduction than general-purpose WavTokenizer and BigCodec. For a single RVQ layer among multiple-codebook acoustic tokens, the reduction effect is also significant. These findings suggest that the effect of BPE is negatively correlated with the information density in the speech tokens: the less information, the more length reduction achieved by BPE.
@@ IV-E Variable Frame Rate Tokens and Unit Discovery
VFR tokens are closely related to acoustic unit discovery. As speech lacks a natural boundary of phonetic units [8], there are much research efforts to find and locate the underlying acoustic units behind speech utterances in an unsupervised manner [164, 165, 166, 167]. This is particularly of interest for low-resource languages. The discovered units can guide the boundary segmentation of VFR tokens. To this end, VFR tokens are interesting not only because they might reduce the necessary bitrate, but also because they can introduce a strong inductive bias that linguistic knowledge is encoded [163].
A recent direction of VFR tokens is to discover acoustic units from an SSL model. Note that deduplicated tokens and acoustic BPE themselves can be regarded as VFR tokens. Sylber [168] and SyllableLM [45] take similar approaches that first locate acoustic boundaries from existing HuBERT models, and then train another HuBERT student with segment-level pooled targets between boundaries. The final HuBERT embeddings undergo the same segment-level pooling and kmeans clustering procedure to produce tokens at a very low frame rate ($\approx 5$Hz) that align well with syllables.
Boundary prediction can be involved to achieve frame rate variability in the training process, where a specific model predicts frame-level boundaries and is trained together with other network modules. The training techniques of such models include reinforcement learning [169], soft pooling [148], and slowness constraint [163]. However, these approaches are rarely adopted in the context of discrete speech tokens today, and there have barely been a VFR acoustic token till now.